The magnetic field B and the magnetic intensity H in a material are found to be 1.6 T and 1000 A/m respectively, Calculate the relative permeability μ^r and the susceptibility 𝒳 of the material.

Grade:11

2 Answers

Kevin Nash
9 years ago
Sol. B = 1.6 T, H = 1000 A/m μ = Permeability of material μ = B/H = 1.6/1000 = 1.6 x 10^-3 μ = μ/ μ base 0 = 1.6x 10^-3 / 4π x 10^-7 = 0.127 x 10^4 ≈ 1.3 x 10^3 μ = μ base 0 (1 + x) ⇒ X = μ/ μ base 0 -1 = μr – 1 = 1.3 × 10^3 – 1 = 1300 – 1 = 1299 ≈ 1.3 × 10^3
Padvi Yogesh
5 years ago
The magnetic field B found the Magnetic intensity H in a material are found to be 1.6 T and 1000A/m respectively.Calculate the relative pearmeability ur and the susceptibility X of the material ?